OpenStudy (anonymous):

I need the area of circle that is 2" in with the DIA of the disk is 8.5"

OpenStudy (amistre64):

what??

OpenStudy (anonymous):

I have a disk at is 8.5 inches in Dia. there is 2 inches of this disk I am trying to figure out the area.

OpenStudy (amistre64):

so if we had a circle and you chopped off an end of it, how much area would it be? like this?

OpenStudy (anonymous):

The 2 inches is the dia too (a complete circle) not a segment.

OpenStudy (amistre64):

like this then? maybe?

OpenStudy (anonymous):

yes something like that. I know the raduis is 4.25

OpenStudy (anonymous):

let say the white part is the 2 "

OpenStudy (amistre64):

ohhh...ok

OpenStudy (amistre64):

diam/2 = 8.5/2 = 4.25 the area of the whole thing then is: (4.25)^2pi now remove the area of the inner circle which has the radius of 2.25 (4.25)^2 pi - (2.25)^2 pi = pi( 4.25^2 - 2.25^2)

OpenStudy (amistre64):

13pi is what I come up with

OpenStudy (anonymous):

Where does the 2.25 come from?

OpenStudy (anonymous):

Thanks I think I figured it out!

OpenStudy (amistre64):

:) I knew you could do it :)